OpenStudy (anonymous):

need help with finding an expression for (15 - 4x) - (10 -x)

OpenStudy (aum):

(15 - 4x) - (10 -x) = 15 - 4x - 10 + x = ? (simplify)

OpenStudy (anonymous):

so it will be (5-3x)

OpenStudy (aum):

yes.